304x0 206 18) The amount of gas that occupies 36.52 L at 68.0 C and 672 mm Hg is A) 127 B) 1.15 6 36.52x0 88 C) 878 D)24.4 E) 12.7 Answer: B Diff. 2 Var: 1 Page Ref Sec. 10.4 LO: 10.4 GO: G4
Answer #1

we have:

P = 672.0 mm Hg

= (672.0/760) atm

= 0.8842 atm

V = 36.52 L

T = 68.0 oC

= (68.0+273) K

= 341 K

find number of moles using:

P * V = n*R*T

0.8842 atm * 36.52 L = n * 0.08206 atm.L/mol.K * 341 K

n = 1.15 mol

Answer: B

Feel free to comment below if you have any doubts or if this answer do not work. I will edit it if you let me know